# GlobalGeoTree: A Multi-Granular Vision-Language Dataset for Global Tree Species Classification

### v2

@@ 5 Ethics, Limitations and Impact
Ethics
GlobalGeoTree is constructed using publicly available data from sources like GBIF [46], Sentinel-2 [47], WorldClim [34], SoilGrids [33], and USGS SRTM [32], all under open licenses (CC0 1.0, CC-BY-4.0, CC-BY-NC-4.0). No personal or sensitive information is included, ensuring ethical use.
Limitations
@@
Despite our efforts to ensure global coverage, the dataset exhibits geographic biases due to uneven species distribution data. Regions with longer histories of biodiversity documentation are overrepresented such as Europe and North America. The ambiguous boundary between trees and shrubs in botanical classification further complicates the dataset, as some samples may represent shrubs rather than trees. Tree species taxonomy is also subject to frequent revisions driven by new genetic evidence, which may misalign dataset labels with updated classifications over time. Although CLIP-based models can identify unseen species to some extent, such taxonomic shifts may still affect model interpretability and evaluation consistency. Additionally, the dataset relies on Sentinel-2 data from 2020, restricting its ability to capture long-term vegetation dynamics or recent disturbances. Future versions incorporating multi-year observations could better account for phenological changes and climate-driven impacts.
Potential impact
GlobalGeoTree holds great potential for advancing forest monitoring, biodiversity conservation, and climate change mitigation. By improving tree species mapping, it can support sustainable forest management, restoration planning, accurate carbon stock estimation, and biodiversity monitoring. However, there is a risk of misuse, such as in deforestation or logging. Responsible use and adherence to conservation principles are essential.
6 Availability and maintenance
@@
The dataset access, pretrained model checkpoints, and all relevant codes are available in our github repository (https://github.com/MUYang99/GlobalGeoTree), which provides comprehensive tools for using these resources. The pretraining dataset GlobalGeoTree-6M and evaluation dataset GlobalGeoTree-10kEval are provided in WebDataset format [48] and hosted on Huggingface (https://huggingface.co/datasets/yann111/GlobalGeoTree). This format enables efficient online data streaming to train models without requiring full dataset downloads, facilitating large-scale machine learning workflows. It also integrates seamlessly with popular deep learning frameworks, improving accessibility and usability for researchers.
@@
We are dedicated to maintaining and enhancing the dataset, addressing issues, and incorporating updates like new data sources or taxonomic revisions in future versions. The code and Huggingface repository will serve as the primary channels for updates and community feedback.
@@ 7 Conclusion and future work
Future work could explore several promising directions. Expanding the GlobalGeoTree with more recent data, additional satellite sensors (e.g., SAR data for structural information), and a broader range of auxiliary variables could enhance its utility. Investigating alternative vision-language model architectures, pretraining strategies, and methods for addressing the long-tail distribution could further improve classification accuracy, especially at the species level. Developing techniques for uncertainty estimation and improving model explainability are also critical areas for future work. Moreover, applying GlobalGeoTree and GeoTreeCLIP to real-world applications in biodiversity monitoring, conservation, and forest management could provide practical support and holds great potential.
Appendix
@@ Appendix A GlobalGeoTree Dataset Statistics
A.1 Basic Statistics
@@
  • Countries/Regions Covered: 221
@@
A.2 Long-tail Distribution Analysis
@@
A.3 Detailed Categorical Statistics
@@
Functional Type (level0)
@@
Taxonomic Family (level1_family)
@@
Taxonomic Genus (level2_genus)
@@
Taxonomic Species (level3_species)
@@
Appendix B Details of Evaluation Subsets
B.1 Overview and Construction
To enable robust benchmarking across various taxonomic diversity scales and species rarity, we constructed three evaluation subsets: GlobalGeoTree-10kEval, GlobalGeoTree-10kEval-300, and GlobalGeoTree-10kEval-900. These subsets were created by first categorizing all species in the GlobalGeoTree dataset into Rare, Common, and Frequent groups based on available sample counts (Section 3.3), then randomly selecting 30, 100, and 300 species per category, respectively. The primary GlobalGeoTree-10kEval benchmark (90 species) is featured in the main paper, while the larger subsets enable assessment of model scalability and performance on increasingly complex tasks. Detailed overviews of each subset’s composition are provided in Table 5, Table 6, and Table 7. The geographical distribution of the two additional evaluation sets is shown in Figure 6(a) and Figure 6(b).
B.2 GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
(a) GlobalGeoTree-10kEval-300
(b) GlobalGeoTree-10kEval-900
@@ B.2 GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
Appendix C Model Performance on GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
The evaluation results on the larger GlobalGeoTree-10kEval-300 (Table 8) and GlobalGeoTree-10kEval-900 (Table 9) subsets align with trends from the primary benchmark GlobalGeoTree-10kEval (Table 3 and 4). GeoTreeCLIP consistently outperforms CLIP and RemoteCLIP across all settings (zero-shot, one-shot, three-shot) and taxonomic levels (Family, Genus, Species). Despite lower absolute accuracies on these challenging subsets, especially GlobalGeoTree-10kEval-900, GeoTreeCLIP maintains a significant performance edge, highlighting the advantages of its domain-specific pretraining and tailored architecture for multimodal tree species classification.
@@ Appendix C Model Performance on GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
Appendix D Comparison with Supervised Learning Paradigm
D.1 Supervised Model Architecture and Training
D.2 Zero-Shot Evaluation of the Supervised Model
Notably, the GlobalGeoTree-6M dataset was designed to retain nearly all tree species categories, as our goal was to pretrain a model capable of classifying the full spectrum of tree species. Consequently, the zero-shot evaluation here can be seen as measuring the model’s “zero-shot” transfer capability on unseen datasets, aligning with the concept of "in-domain" zero-shot classification defined in [25].
Table 10: Zero-shot evaluation on GlobalGeoTree-10kEval. SupervisedGeoTree is evaluated on classes within its training vocabulary. Results are mean accuracy (%) $\pm$ standard deviation (%) over 5 runs.
Taxon. CLIP RemoteCLIP SupervisedGeoTree GeoTreeCLIP
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
Family 10.80 $\pm$ 0.03 25.32 $\pm$ 0.05 1.11 $\pm$ 0.01 10.55 $\pm$ 0.04 9.55 $\pm$ 0.10 27.41 $\pm$ 0.07 20.99 $\pm$ 0.28 56.88 $\pm$ 0.42
Genus 1.09 $\pm$ 0.01 9.34 $\pm$ 0.01 1.11 $\pm$ 0.01 6.25 $\pm$ 0.02 1.19 $\pm$ 0.09 8.18 $\pm$ 0.14 18.39 $\pm$ 0.26 50.98 $\pm$ 0.41
Species 1.09 $\pm$ 0.01 7.02 $\pm$ 0.02 1.11 $\pm$ 0.01 6.25 $\pm$ 0.02 0.00 $\pm$ 0.00 0.28 $\pm$ 0.02 16.71 $\pm$ 0.25 47.52 $\pm$ 0.37
In stark contrast, our proposed GeoTreeCLIP model demonstrates substantial improvements over all other baselines across every taxonomic level. At the family level, GeoTreeCLIP achieves a top-1 accuracy of 20.99%, more than doubling the performance of CLIP and significantly outperforming SupervisedGeoTree. This advantage is even more pronounced at the finer-grained levels: GeoTreeCLIP obtains 18.39% top-1 accuracy for genus and 16.71% for species identification. These results strongly indicate the power of contrastive vision-language learning to align visual features with rich, hierarchical taxonomic text labels. This capability allows GeoTreeCLIP to learn more nuanced and generalizable representations, leading to its superior performance in this challenging zero-shot evaluation. The significant gap, especially compared to SupervisedGeoTree at the species level, highlights the efficacy of the contrastive learning paradigm for handling large, structured label spaces.
Appendix E Qualitative Analysis: t-SNE Feature Embeddings Visualization
To qualitatively assess and compare the learned feature representations from different modeling paradigms, we performed t-SNE visualizations [49]. We used zero-shot image embeddings (or the final fused visual-auxiliary features for SupervisedGeoTree before the classification heads) extracted from a subset of the GlobalGeoTree-10kEval-300 dataset. This analysis includes our proposed GeoTreeCLIP, the original CLIP pretrained by OpenAI, and the SupervisedGeoTree model. Given the large number of classes and the hierarchical nature of the labels, we adopted a selective visualization strategy: first examining embeddings at the family level for five randomly selected families (Anacardiaceae, Berberidaceae, Cactaceae, Hernandiaceae, and Rosaceae); then focusing on five genera within the Rosaceae family (Prunus, Pyrus, Rhodotypos, Rubus, and Spiraea); and finally, visualizing five species within the Prunus genus (Prunus avium, Prunus caroliniana, Prunus ilicifolia, Prunus laurocerasus, and Prunus obtusata).
The comparative t-SNE visualizations are presented in Figure 7. Across all three taxonomic levels (Family, Genus, and Species, shown as columns), GeoTreeCLIP (top row) consistently demonstrates the most effective separation and formation of distinct clusters. At the family level (left column), GeoTreeCLIP clearly distinguishes between the selected families. The original CLIP (middle row) exhibits considerable overlap, particularly for the diverse Rosaceae family. SupervisedGeoTree (bottom row) shows some separation but less defined clusters compared to GeoTreeCLIP, with Rosaceae still forming a very broad distribution.
Figure 7: t-SNE visualization of feature embeddings from GlobalGeoTree-10kEval-300 at different taxonomic levels, comparing GeoTreeCLIP (top row), original CLIP (middle row), and SupervisedGeoTree (bottom row). Columns from left to right represent visualizations at the Family level (selected: Anacardiaceae, Berberidaceae, Cactaceae, Hernandiaceae, Rosaceae), Genus level (selected within Rosaceae: Prunus, Pyrus, Rhodotypos, Rubus, Spiraea), and Species level (selected within Prunus: Prunus avium, Prunus caroliniana, Prunus ilicifolia, Prunus laurocerasus, Prunus obtusata).
This pattern of superior clustering by GeoTreeCLIP continues at the genus level within Rosaceae (middle column). GeoTreeCLIP forms relatively distinct groups for genera like Prunus, Pyrus, and Rhodotypos. Both CLIP and SupervisedGeoTree struggle more, with CLIP showing a highly condensed and overlapping structure, while SupervisedGeoTree offers some separation but with less clarity than GeoTreeCLIP. The most striking difference is observed at the species level within the Prunus genus (right column). GeoTreeCLIP achieves remarkable separation, forming visually distinct clusters for each of the five Prunus species. In contrast, both original CLIP and SupervisedGeoTree largely fail to differentiate these closely related species, with their embeddings heavily intermingled.
Appendix F Additional Baselines Evaluated on GlobalGeoTree-10kEval
@@
To further contextualize GeoTreeCLIP’s performance, we extended our zero-shot evaluation on the GlobalGeoTree-10kEval subset to include two additional publicly available vision-language models: SkyCLIP-50 [25] and CLIP-laion-RS, a CLIP model pretrained on the remote sensing subset of LAION-2B [50]. These models were evaluated on the GlobalGeoTree-10kEval subset under the same zero-shot protocol used for original CLIP and RemoteCLIP (features extracted from individual monthly images, probabilities averaged). For ease of comparison, their performance alongside our GeoTreeCLIP is presented in Table 11.
@@
The results in Table 11 show that both SkyCLIP-50 and CLIP-laion-RS, despite their pretraining on remote sensing imagery, achieve zero-shot accuracies on GlobalGeoTree-10kEval that are substantially lower than our GeoTreeCLIP. For instance, at the species level, SkyCLIP-50 obtains a top-1 accuracy of 1.10% and CLIP-LAION-RS achieves 1.12%, in contrast to GeoTreeCLIP’s 16.71%. As indicated in the main text for similar baseline models (original CLIP, RemoteCLIP), such performance can partly be attributed to their design, which is often optimized for RGB data and lacks effective handling of time-series information or small-patch classification crucial for tree species identification. These limitations further emphasize the value of the GlobalGeoTree benchmark and the effectiveness of our tailored GeoTreeCLIP approach in advancing global tree species classification research.

### v3

@@ 7 Ethics, Limitations and Impact
7.1 Ethics
Our data collection process was designed to adhere to ethical standards for data reuse. All geolocated occurrence records sourced from GBIF (Lane and Edwards, 2007) were filtered to include only those with permissive licenses (e.g., CC0, CC-BY) explicitly chosen by the original data contributors. We also respected the privacy safeguards of the source platforms; for instance, the coordinates of sensitive or threatened species are often automatically obscured by platforms like iNaturalist, and our dataset preserves these privacy-preserving modifications.
The Sentinel-2 L2A data (Spoto et al., 2012) were accessed via Google Earth Engine. The auxiliary bioclimatic, geographic, and soil data were sourced from WorldClim (Fick and Hijmans, 2017), USGS SRTM (Jarvis et al., 2008), and SoilGrids (Poggio et al., 2021), respectively, all of which are publicly available.
7.2 Limitations
@@
While GlobalGeoTree represents one of the largest and most comprehensive datasets of its kind, users should be aware of its inherent limitations. As with large part of datasets derived from citizen-science observations, it contains geographic and taxonomic biases. Data coverage is higher in regions with active observer communities (e.g., North America, Europe) and is skewed towards more common or easily identifiable species, resulting in a long-tail distribution. Furthermore, since we assimilated available official inventory data (e.g., NFIs) into the training set to maximize coverage, independent external validation remains challenging, and shared observation biases may influence evaluation metrics. The ambiguous boundary between trees and shrubs in botanical classification further complicates the dataset, as some samples may represent shrubs rather than trees. Tree species taxonomy is also subject to frequent revisions driven by new genetic evidence, which may misalign dataset labels with updated classifications over time. Although CLIP-based models can identify unseen species to some extent, such taxonomic shifts may still affect model interpretability and evaluation consistency. Additionally, the dataset relies on Sentinel-2 data from 2020, restricting its ability to capture long-term vegetation dynamics or recent disturbances. Future versions incorporating multi-year observations could better account for phenological changes and climate-driven impacts.
7.3 Potential impact and Operational Deployment
GlobalGeoTree holds significant potential for advancing forest monitoring, biodiversity conservation, and climate change mitigation. Beyond benchmarking, the dataset and GeoTreeCLIP model support practical operational workflows. For large-scale mapping, users can employ a sliding-window inference approach on Sentinel-2 tiles to generate pixel-wise species maps. To handle prediction uncertainty in diverse ecosystems, we recommend a hierarchical inference strategy, where the model defaults to genus or family-level classifications if species-level confidence falls below a threshold. Furthermore, to address regional domain shifts, the pretrained model serves as a robust foundation that can be efficiently fine-tuned with small sets of local field data. This "local adaptation" workflow allows practitioners to leverage global knowledge while tailoring predictions to specific forest management needs.
8 Code and Data Availability
@@
The GlobalGeoTree dataset is openly available under the Creative Commons Attribution 4.0 International (CC-BY 4.0) license. The complete dataset is archived on Hugging Face at https://huggingface.co/datasets/yann111/GlobalGeoTree under DOI https://doi.org/10.15468/dd.9qxqyy (Mu et al., 2025b). The pretraining and evaluation datasets are provided in WebDataset format (Aizman et al., 2019). This format enables efficient online data streaming to train models without requiring full dataset downloads, facilitating large-scale machine learning workflows. It also integrates seamlessly with popular deep learning frameworks, improving accessibility and usability for researchers.
@@
All code is available in our github repository https://github.com/MUYang99/GlobalGeoTree under the Apache License 2.0, which provides comprehensive tools for using these resources. We are dedicated to maintaining and enhancing the dataset, addressing issues, and incorporating updates like new data sources or taxonomic revisions in future versions. The code and Huggingface repository will serve as the primary channels for updates and community feedback.
@@ 9 Conclusion and future work
Future work could explore several promising directions. Expanding the GlobalGeoTree with more recent data, additional satellite sensors (e.g., SAR data for structural information), and a broader range of auxiliary variables could enhance its utility. Specifically, incorporating global structural datasets, such as GEDI LiDAR metrics or Meta’s 1-m Canopy Height Map, could provide critical vertical information to distinguish between ecologically similar tree and shrub species, a limitation of current spectral-only approaches. Investigating alternative vision-language model architectures, pretraining strategies, and methods for addressing the long-tail distribution could further improve classification accuracy, especially at the species level. Developing techniques for uncertainty estimation and improving model explainability are also critical areas for future work. Moreover, applying GlobalGeoTree and GeoTreeCLIP to real-world applications in biodiversity monitoring, conservation, and forest management could provide practical support and holds great potential.
@@ Appendix A GlobalGeoTree Dataset Statistics
A.1 Data Records and Taxonomic Catalog
To provide full transparency and facilitate detailed exploration of the dataset, we have made two key files publicly available in our Hugging Face repository (https://huggingface.co/datasets/yann111/GlobalGeoTree/files).
  • Complete Taxonomic Catalog (Tree_species_catalog.csv): This file contains the complete list of all 275 families, 2,734 genera, and 21,001 species included in GlobalGeoTree, along with their corresponding unique species keys used in the GBIF database. This catalog serves as a comprehensive reference for the taxonomic scope of our dataset.
  • Full Georeferenced Records (GlobalGeoTree.csv): This file contains all 6.3 million geolocated records. Each row includes the coordinates, hierarchical taxonomic labels (from functional type to species), the original data source (e.g., iNaturalist Research-grade Observations), and other metadata as detailed in Table 3. This allows for complete traceability and enables users to perform custom geographical or taxonomic subsetting.
A.2 Basic Statistics
@@
  • Countries/Regions [^1] The term ”Countries/Regions” is used to align with global data standards (e.g., from GBIF) and accurately represents both sovereign states and other geographic entities such as overseas territories or dependencies, ensuring comprehensive coverage. Covered: 221
@@
A.3 Long-tail Distribution Analysis
@@
A.4 Detailed Categorical Statistics
@@ A.4.1 Location (location)
A.4.2 Functional Type (level0)
@@
A.4.3 Taxonomic Family (level1_family)
@@
A.4.4 Taxonomic Genus (level2_genus)
@@
A.4.5 Taxonomic Species (level3_species)
@@
A.5 Representative Species Profile
Table 10 provides a structured summary of representative species from the GlobalGeoTree dataset, highlighting the diversity across different functional types and geographic regions.
Table 10: Structured summary of representative tree species in GlobalGeoTree, spanning major functional types and biogeographical realms.
Functional Type Family Genus Species Dominant Region Ecological Context
Evergreen Needleleaf Pinaceae Picea Picea abies Europe Dominant boreal forest conifer
Evergreen Needleleaf Pinaceae Pseudotsuga Pseudotsuga menziesii North America Key timber species of Pacific Northwest
Deciduous Broadleaf Fagaceae Quercus Quercus robur Europe Keystone species of temperate woodlands
Deciduous Broadleaf Sapindaceae Acer Acer saccharum North America Major component of northern hardwood forests
Evergreen Broadleaf Myrtaceae Eucalyptus Eucalyptus globulus Oceania Fast-growing, widely distributed in Australia
Evergreen Broadleaf Urticaceae Cecropia Cecropia peltata South America Pioneer species in neotropical rainforests
Deciduous Needleleaf Pinaceae Larix Larix decidua Europe High-altitude conifer of the Alps
Appendix B Auxiliary Environmental Variables
The 27 auxiliary variables used in GlobalGeoTree were sourced from three publicly available datasets, accessible via Google Earth Engine (GEE):
  • WorldClim V1 (Bioclimatic Variables): GEE Dataset ID: WORLDCLIM/V1_BIO
  • SoilGrids 250m v2.0 (Soil Variables): GEE Dataset ID: ISRIC/soilgrids/v02
  • USGS SRTM GL1 v003 (Geographic Variables): GEE Dataset ID: USGS/SRTMGL1_003
Table LABEL:tab:bioclim_vars provides a detailed description of a subset of these variables.
Table 11: Description of the 19 bioclimatic variables from WorldClim (BIO1-BIO19) used in the GlobalGeoTree dataset.
Variable Units Scale Factor Description
[…]
bio03 % 1 Isothermality (BIO2/BIO7 $\times$ 100)
[…]
bio12 mm 1 Annual Precipitation
bio13 mm 1 Precipitation of Wettest Month
bio14 mm 1 Precipitation of Driest Month
bio15 Coeff. of Variation 1 Precipitation Seasonality
bio16 mm 1 Precipitation of Wettest Quarter
bio17 mm 1 Precipitation of Driest Quarter
bio18 mm 1 Precipitation of Warmest Quarter
bio19 mm 1 Precipitation of Coldest Quarter
Appendix C Details of Evaluation Subsets
C.1 Overview and Construction
To enable robust benchmarking across various taxonomic diversity scales and species rarity, we constructed three evaluation subsets: GlobalGeoTree-10kEval, GlobalGeoTree-10kEval-300, and GlobalGeoTree-10kEval-900. These subsets were created by first categorizing all species in the GlobalGeoTree dataset into Rare, Common, and Frequent groups based on available sample counts (Sect. 3.4), then randomly selecting 30, 100, and 300 species per category, respectively. The primary GlobalGeoTree-10kEval benchmark (90 species) is featured in the main paper, while the larger subsets enable assessment of model scalability and performance on increasingly complex tasks. Detailed overviews of each subset’s composition are provided in Table 12, Table 13, and Table 14. The geographical distribution of the two additional evaluation sets is shown in Figure 10a and Figure 10b.
C.2 GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
@@
(a) GlobalGeoTree-10kEval-300
(b) GlobalGeoTree-10kEval-900
Appendix D Model Performance on GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
The evaluation results on the larger GlobalGeoTree-10kEval-300 (Table 15) and GlobalGeoTree-10kEval-900 (Table 16) subsets align with trends from the primary benchmark GlobalGeoTree-10kEval (Table 4 and 7). GeoTreeCLIP consistently outperforms CLIP and RemoteCLIP across all settings (zero-shot, one-shot, three-shot) and taxonomic levels (Family, Genus, Species). Despite lower absolute accuracies on these challenging subsets, especially GlobalGeoTree-10kEval-900, GeoTreeCLIP maintains a significant performance edge, highlighting the advantages of its domain-specific pretraining and tailored architecture for multimodal tree species classification.
@@ Appendix D Model Performance on GlobalGeoTree-10kEval-300 and GlobalGeoTree-10kEval-900
Appendix E Additional Baselines Evaluated on GlobalGeoTree-10kEval
@@
To further contextualize GeoTreeCLIP’s performance, we extended our zero-shot evaluation on the GlobalGeoTree-10kEval subset to include two additional publicly available vision-language models: SkyCLIP-50 (Wang et al., 2024) and CLIP-laion-RS, a CLIP model pretrained on the remote sensing subset of LAION-2B (Schuhmann et al., 2022). These models were evaluated on the GlobalGeoTree-10kEval subset under the same zero-shot protocol used for CLIP and RemoteCLIP (features extracted from individual monthly images, probabilities averaged). For ease of comparison, their performance alongside our GeoTreeCLIP is presented in Table 17.
@@
The results in Table 17 show that both SkyCLIP-50 and CLIP-laion-RS, despite their pretraining on remote sensing imagery, achieve zero-shot accuracies on GlobalGeoTree-10kEval that are substantially lower than our GeoTreeCLIP. For instance, at the species level, SkyCLIP-50 obtains a top-1 accuracy of 1.10% and CLIP-LAION-RS achieves 1.12%, in contrast to GeoTreeCLIP’s 16.71%. As indicated in the main text for similar baseline models (CLIP, RemoteCLIP), such performance can partly be attributed to their design, which is often optimized for RGB data and lacks effective handling of time-series information or small-patch classification crucial for tree species identification. These limitations further emphasize the value of the GlobalGeoTree benchmark and the effectiveness of our tailored GeoTreeCLIP approach in advancing global tree species classification research.
@@
[…]
Conceptualization: Y.M., Z.X., S.M., X.Z.; methodology: Y.M., Z.X., Y.W., S.M., X.Z.; software: Y.M., Z.X., Y.W., X.Z.; results validation: Y.M., Z.X., Y.W., S.M., F.E., H.K., M.K., X.Z.; analysis: Y.M., Z.X., X.Z.; data collection: Y.M., Y.W.; writing, and original draft preparation: Y.M.; paper revision: Y.M., Z.X., Y.W., S.M., F.E., H.K., M.K., X.Z.; visualization: Y.M., Z.X.; supervision: S.M., X.Z.; funding acquisition: X.Z.; project administration: X.Z.; resources: X.Z.
[…]
none
[…]
none
Acknowledgements.
The work is jointly supported by the German Federal Ministry for Economic Affairs and Climate Action in the framework of the ”national center of excellence ML4Earth” (grant number: 50EE2201C), by the European Union’s Horizon Europe research and innovations actions programme in the framework of the ”Multi-source and Multi-scale Earth observation and Novel Machine Learning Methods for Mineral Exploration and Mine Site Monitoring (MultiMiner)” (grant number: 101091374) and “ThinkingEarth—Copernicus Foundation Models for a Thinking Earth” (Grant Agreement No. 101130544), by the German Federal Ministry for the Environment, Nature Conservation, Nuclear Safety and Consumer Protection (BMUV) based on a resolution of the German Bundestag (grant number: 67KI32002B; Acronym: EKAPEx), by the European Research Council (ERC) (grant agreement No. [ERC-2022-PoC-101101093], Acronym: EO4FoodSecurity) by the Excellence Strategy of the Federal Government and the Länder through the TUM Innovation Network EarthCare and by Munich Center for Machine Learning.